OpenStudy (anonymous):

-8=1/3x+x please solve and show the steps thank you

OpenStudy (anonymous):

-8=1/3x+x 4x/3= -8 4x=-24 x=-6

OpenStudy (anonymous):

how did u get 4x/3= -8

OpenStudy (anonymous):

1/3x+x=-8 here LCM is 3 , therefor (1x+3x)/3= -8 ie 4x/3=-8 got it?

OpenStudy (anonymous):

please tell me :)

OpenStudy (anonymous):

which part you don't understand?